How many electrons are present in a wire if a 1.5 A current passes through the conductor in 2 minutes?

1
Expert's answer
2022-03-29T12:27:52-0400
"N=\frac{q}{e}=\frac{It}{e}"

"N=\frac{1.5*2*60}{1.6*10^{-19}}=1.1*10^{21}"